a 150g ball at the end of a string is revolving uniformly in a horizontal circle of radius 0.6 m .the ball makes 2 revolutions in a second what is it’s centripal acceleration


1
Expert's answer
2021-06-13T07:09:16-0400

"a=v^2/r=\omega^2r"


"2\ (rev/s)=4\pi\ (rad/s)" . So,


"a=\omega^2r=(4\pi)^2\cdot0.6=94.75\ (m/s^2)" . Answer